# Epoch-based Optimistic Concurrency Control in Geo-replicated Databases

### v1

@@ 3.2. Optimistic Local Execution
3: var temp_state : DatabaseState $\triangleright$ Local uncommitted changes since the last consistent snapshot
@@
6:   if accessed items are last updated by txntxns in the same epoch then $\triangleright$ Dependencies tracking for conflict detection
7:    add the $txn.tid$s to $txn.dependencies$
8:   else if accessed items lasted updated in previous epochs then
@@
During the read phase, transactions access a temporary state containing uncommitted local changes applied atop the most recent consistent snapshot. This snapshot reflects the database state at the conclusion of the previous epoch. The temporary state is reset to the new snapshot at the start of every new epoch.
Data items written to the temporary state are tagged with the tidtid of the updating transaction. This mechanism is used to track local dependent transactions, that is, transactions that read data items updated by other transactions within the same epoch. These dependencies are critical for cascading aborts: if a writer transaction ($T_{write}$) is aborted, any dependent reader ($T_{read}$) must also be aborted, as its read-set has become invalid. Conversely, when a transaction reads an unmodified item directly from the consistent snapshot, it records the item’s version, denoted by the epoch ID (eideid) of its last update. This eideid is utilized during the global commit phase (§3.4) to detect stale reads (i.e., cases where the local snapshot was superseded by concurrent remote updates).
Following execution, the transaction’s input, read-set, and write-set are appended to the current batch. Upon reaching capacity or after a timeout, the batch is asynchronously broadcast to all replicas ($BroadcastBatchAsync()$). Simultaneously, these batches are persisted into a linear, per-replica local log, ensuring availability for the subsequent commit protocol.
@@
Figure 3 illustrates the organization of locally executed transactions within the log of batches, where arrows denote data dependencies. In this example, $7$ transactions are distributed across two batches. Transactions $tx1$, $tx2$, $tx4$, and $tx7$ are independent; they operate on disjoint keys and read directly from the consistent snapshot. In contrast, $tx3$, $tx5$, and $tx6$ are dependent transactions, reading values written by earlier, uncommitted local operations.
For instance, $tx3$ depends on $tx1$ and $tx2$ (reading data item $A$ and writing to $B$), which implies a strict execution order where $tx3$ succeeds its predecessors. Similarly, $tx5$ exhibits dependencies on earlier writes. Transaction $tx6$, reading both $A$ and $B$, depends on both $tx5$ and $tx3$. The blue circles encircle these dependency groups, which we designate as Transaction Chains (TxnC). These chains serve as the atomic units for conflict detection during the global validation phase, as discussed in detail in Section 3.5.
@@ 3.3. Asynchronous Replication
11:    for $i\leftarrow PoA\_heads[self.rid]+1$ to $ack.bid$ do $\triangleright$ If a batch and its predecessors are confirmed available, broadcast PoA until the latest available one.
@@
Figure 4. Replica A’s view of the logs: Brown means committed, blue means available and has PoA, green means broadcast but without PoA, and white means received but not available. Red dotted line is the coordinator’s consistent cut.
Upon receiving a batch from a peer, a replica appends the entry to the corresponding source replica’s log, marks it as broadcast, and sends an acknowledgment ($ReceivedBatch()$). The source replica aggregates these acknowledgments in $receivedAcknowledgement()$. After securing a quorum of $f+1$ acknowledgments, the batch transitions to the available state (Line 10). At this point, the source replica broadcasts a Proof of Availability (PoA), but only when all preceding batches in the sequence have already achieved available status and issued their own PoAs. This constraint ensures that availability is certified only for contiguous log prefixes. The PoA serves as a guarantee that the batch is persisted on at least one correct replica, satisfying the $f+1$ availability requirement.
The PoA generation process (Line 11) iteratively scans the log from the current PoA_headPoA\_head (the index of the last batch with a transmitted PoA) up to the newly available batch. To ensure log contiguity, a PoA is broadcast for a batch only if all its predecessors are also marked available and the PoA_headPoA\_head is advanced. Upon receiving a PoA ($receivedPoA()$), peer replicas update the corresponding batch’s state to PoA_sentPoA\_sent. Crucially, if a replica receives a PoA for a batch that is absent from its local log, it broadcasts a fetch request to retrieve the missing batch.
@@ 3.4. Global Epoch Commit
2: function ProposeConsensus $\triangleright$ At the coordinator
@@
4:    ProposeConsensus(PoA_headsPoA\_heads)
@@
7:   for all rids do
8:    for $i\leftarrow last\_commited[rid]$ to cut_bids[rid]cut\_bids[rid] do
@@
Periodically (every epoch), a designated coordinator replica (typically the Raft leader) initiates a consensus round triggered by a local timer ($ProposeConsensus()$). The coordinator proposes a ”consistent cut” of the global log, consisting of the vector of batch indices currently marked as PoA_sent in its local view (PoA_headsPoA\_heads). Standard consensus protocols, such as Raft (diegoRaft2014), are used for this proposal. Unlike systems that depend on synchronized physical clocks or centralized sequencers for ordering (thomsonCalvinFastDistributed2012; fanOceanVistaGossipbased2019; luEpochbasedCommitReplication2021; zhougeogauss2023), Minerva reaches agreement directly on the set of available batches. This approach establishes a consistent global checkpoint without assuming tight clock synchronization or imposing hard real-time constraints.
@@ 3.5. Validation and Deterministic Re-execution
Algorithm 4 Validation
1: var current_eid : int $\triangleright$ ID of the current epoch
@@
After replicas have collected the batches_to_commitbatches\_to\_commit, replicas independently complete the validation and re-execution phase without further coordination.
We define two types of invalid transactions. The first are stale transactions: those that read a data version from a consistent snapshot that has been superseded by a committed update prior to the current epoch. We use the epoch ID (eideid) as write versioning to detect stale reads (Line 8 in Algorithm 5). In addition, we enforce cascading invalidation: any transaction that depends on a previously invalidated transaction, even across epochs, is recursively marked as stale. The second type is conflicting transactions, which includes read-write or write-write conflicts in the current epoch.
@@
Following conflict resolution, the write-sets of all valid transactions are committed to durable storage. A key aspect of our algorithm is to deterministically re-execute the conflicting transactions instead of aborting them. To ensure serializability, stale and conflicting transactions are first arranged into a total order determined by their (tidtid) and the predefined priority of their source replica. This queue is then executed using a deterministic locking manager, derived from Calvin (thomsonCalvinFastDistributed2012), to enforce serializable execution according to the sorted order while maintaining concurrency. Deterministic locking require prior knowledge of the read and write sets, which are available in our design as they are tracked during the initial OCC phase. Once the results of the re-executed transactions are persisted, the system responds to the clients with the results.
@@ 3.6. Detecting Conflicting Transactions
Given the batch of transactions, we must efficiently identify write-write and read-write conflicts among them to ensure serializability (bernsteinconcurrency1981; ceriOCC1982). The conflict detection process must be deterministic, ensuring that every replica independently reaches the exact same commit decision; optimal, minimizing the number of re-executions; and computationally efficient, ensuring that the validation logic does not become a bottleneck in the commit critical path.
@@
25:   $\triangleright$ Solve Maximum Weight Independent Set to find valid txns
26:   var $valid\_set\leftarrow$ SolveMWIS(CGCG)
27:   var $conflicting\_set\leftarrow CG.vertices\setminus valid\_set$
28:   return $(conflicting\_set\cup stale\_txns,\ valid\_set)$
Figure 5. Conflict Graph of 6 transaction chains, arrows mean read-write or write-write conflicts, and red box shows the conflicting transaction chains (TxnCTxnC) that are invalidated and to be re-executed.
Algorithm 5 shows the approach. We first construct the conflict graph ($CreateConflictGraph()$) through an iterative scan of all write-sets, linking transaction chains that have read-write or write-write contention on shared data items. Subsequently, the resulting graph is processed by the MWIS solver to identify the optimal subset of non-conflicting transactions. The chains that are excluded from the independent set are designated as conflicting. To ensure global consistency in cases where multiple optimal solutions exist, the solver may employ a deterministic tie-breaking rule, such as prioritizing the transaction chain with the lowest starting tidtid.
An example is shown in Figure 5 where $txnc2$ represents a heavy chain containing 5 transactions, while all other nodes ($txnc1$, txnc36txnc3-6) represent chains with a single transaction each. $txnc1$, $txnc2$, and $txnc3$ are in conflict because they may have all accessed the same data item. However, the MWIS solver should remove both $txnc1$ and $txnc3$ because $txnc2$ has the highest weight. For $txnc4$, 55, and $6$, a possible scenario is that there are two data items, $a$ and $b$, $txnc4$ writes to $a$, $txnc5$ writes to both $a$ and $b$, and $txnc6$ reads from $b$. In this case, removing $txnc5$ is optimal.
Another issue to consider is the cross-epoch and cross-batch dependent transactions. For cross-epoch dependency, that is, a transaction depends on other transactions, but they are not committed in the same epoch. In this case, the transaction in the second batch must be derived from the previous consistent snapshot, so it is treated the same as stale transactions. For dependencies across batches but still within the same epoch, they are placed in the same global batch, and the link between the transaction chains across batches can be used to merge them into a single chain when constructing the conflict graph.
@@ 3.7. Correctness
We show Minerva guarantees One-Copy Serializability (1SR) (bernsteinconcurrency1981) under crash failures. That is, all replicas observe the same serializable execution order of transactions. A transaction is defined as a series of read, write or delete operations, each operation only accesses a single data item. We do not assume any range queries or transactions with unbounded read/write sets, which can be supported with additional mechanisms such as index or partition versioning, or client-side lock (levandoskiMultiVersionRangeConcurrency2015).
@@
Because the coordinator only proposes the consistent cut with indices of batches that have received PoA, and a replica only broadcasts a PoA for a batch at index $i$ if all predecessor batches (indices <i<i) that are already marked as available and have sent their PoAs (Line 11 of Algorithm 2). By induction, if the coordinator observes a PoA for index $j$ (the cut index), then it is guaranteed that PoAs (and thus availability) exist for all indices 0,1,,j10,1,...,j-1.
The consensus (Raft) guarantees that consistent cut indices are agreed upon by all correct replicas in the same order. Since any batches with indices less than the cut index are guaranteed to be available by Lemma 3.1, all the batches are guaranteed to be available. Because the batches are uniquely identified by a monotonically increasing index, the order of the batches will be observed in the same order by all correct replicas.
@@
At each epoch, the same set of transactions are ordered to be executed in a serializable manner among all correct replicas.
@@
By Lemma 3.2, all correct replicas observe the same set of batches to commit, thus the same set of all transactions $S_{all}$. The $CreateConflictGraph()$ function in Algorithm 5 is deterministic, as it detects all readwrite and writewrite conflicts among transactions (chains). The MWIS solver is also deterministic and returns the same optimal set of transactions $S_{opt}$ without readwrite and writewrite conflict at each replica. Thus, there is no dependency among these transactions, and they can be considered serializable (thomasianDistributedOptimisticConcurrency1998).
The remaining transactions $S_{re}$ are re-executed using a deterministic locking mechanism. By the definition of deterministic locking (thomsonCalvinFastDistributed2012), $S_{re}$ is also deterministic serialized among all replicas.
Since $S_{re}$ always executed after $S_{opt}$, and $S_{re}+S_{opt}=S_{all}$, the execution order of $S_{all}$ is serializable. ∎
@@
Base Case: Let $E_{0}$ be the initial epoch; it has no transaction and is trivially serializable.
@@
Induction Step: By the induction hypothesis, upon $E_{k}$, all replicas must start in the same state (the consistent snapshot). Thus, all transactions in epoch $E_{k}$ are validated/committed against the consistent snapshot.
By Lemma 3.3, the commit procedure produces an execution order $S_{k}$ of all transactions in $E_{k}$ equivalent to a serial execution at all replicas, such that $E_{k}=E_{k-1}+S_{k}$.
Since the state at the end of epoch $E_{k-1}$ is consistent among all replicas, and $S_{e}$ is serializable, the state at the end of epoch $E_{k}$ is also consistent among all replicas. Thus, $\sum_{i=1}^{k}S_{i}$ is $S_{ser}$. ∎
@@ 4. Implementation
We have implemented Minerva (minerva) as a distributed transactional key-value store. The system is developed in C# using .NET 9.0. For the network layer, we employ raw TCP sockets to handle all inter-replica and client interactions. We utilize MemoryPack (memorypack), a high-performance binary serializer for the communication protocol.
@@
Our deterministic re-execution engine implements a deterministic locking scheduler derived from Calvin (thomsonCalvinFastDistributed2012). Transactions are first assigned with a global order based on a tuple of the source replica’s static priority and the transaction’s original ID (tidtid). A centralized lock manager enforces this schedule using per-key wait queues. Since the read and write sets are already known from the initial OCC attempt, lock acquisition is deterministic and deadlock-free. The lock manager maintains a dependency counter for each pending transaction; as active transactions complete and release their locks, the manager processes the relevant queues. When a transaction’s dependency counter reaches zero, it is scheduled for execution on a dedicated worker thread.
@@
As established in §3.5, we first need to identify all transaction chains within batches from the same replica. Using the local dependencies that are tracked during the OCC phase, the task is to find all connected components within the local dependency graph (distinct from the global conflict graph). This graph treats each transaction as a vertex, with edges linking transactions that share a local dependency. Then, a Depth-First Search traversal partitions the graph into disjoint sets, which produces the transaction chains.
Next, we use an open-source C++ graph library (Solver) that provides both exact and approximate solvers for the MWIS problem. For the exact solution, the Integer Linear Programming optimization method (wolseyIntegerProgramming2020) is used. Although this approach incurs higher computational latency, it guarantees the identification of the optimal independent set and minimizes the volume of transactions to re-execute. The approximate algorithm is based on a popular greedy heuristic (sakaigreedymwis2003), which prioritizes decision speed over optimality. The performance trade-offs between these two approaches are analyzed in Section 5.5.
@@
Minerva eliminates the need for the blocking Two-Phase Commit (2PC) protocol to ensure atomicity within the replication cluster because of the replication protocol and deterministic validation. Replicas can thus independently converge on the same commit or abort decisions for every transaction without additional coordination.
The Raft cluster is also used as the configuration manager. The Raft leader assumes the role of the coordinator replica, which is responsible for initiating the periodic global commit process. This also effectively handles coordinator failures, as the Raft protocol guarantees that a new leader will be elected to resume the commit process. We use the open-source Raft implementation from the .Next library (DotNext).
Finally, we propose the following potential recovery mechanisms to handle replica failures. While failed replicas can recover by fetching the missing batches from other replicas based on the global log, the global log should be persisted on the disk with a write-ahead log, allowing the recovered replica to only gather the missing batches without the need to re-execute all transactions within the epoch. Furthermore, storage nodes can employ a localized le primary-backup mechanism with asynchronous backup. Since transactions are always written to the database snapshot first, we do not need to consider the consistency problem with the backups by using epoch as versioning.
@@
We implement an adaptive high-contention mode to reduce performance impact under heavily contended workloads. This mode is triggered when the percentage of re-executed transactions exceeds a configurable threshold (defaulting to $50\%$) for a sustained duration (default $200$ ms). In this mode, replicas suspend the optimistic execution phase entirely. Instead, incoming transactions are statically probed to extract their read and write sets and are immediately forwarded to the current batch for replication. During the global commit phase, these transactions bypass the conflict detection phase and are automatically routed to the deterministic execution queue.
@@
To prevent unbounded storage growth, the system implements a garbage collection mechanism that prunes obsolete log entries. Since the global log acts as the source of truth, a batch can be safely discarded only after it has been fully processed by the entire cluster. To track this, replicas periodically broadcast their progress (latest committed epoch). Each node maintains a local view of the cluster’s ”low watermark,” which is the highest epoch index that has been universally committed. Once this watermark advances beyond epoch $i$, replicas safely truncate the log up to that point. In addition, the local temp_statetemp\_state is cleared at each epoch boundary, effectively garbage collecting the invalid transactions.

### v2

@@ 3.2. Optimistic Local Execution
3: var temp_state : DatabaseState $\leftarrow$ Snapshot of the previous epoch
@@
6:   if accessed items are last updated by txntxns in the same epoch then
7:    add the $txn.tid$s to $txn.dependencies$ $\triangleright$ Dependencies tracking for conflict detection
8:   else if accessed items last updated in previous epochs then
@@
Transactions access a temporary state containing uncommitted local changes applied atop the most recent consistent snapshot during read. This snapshot reflects the database state at the conclusion of the previous epoch. The temporary state is reset to the new snapshot at the start of every new epoch.
Data items buffered at the temporary state are tagged with the tidtid of the updating transaction. This mechanism is used to track local dependent transactions, that is, transactions that read data items updated by other transactions within the same epoch. These dependencies are critical for cascading aborts: if a writer transaction ($T_{write}$) is aborted, any dependent reader ($T_{read}$) must also be aborted, as its read-set has become invalid. Conversely, when a transaction reads an unmodified item directly from the consistent snapshot, it records the item’s version, denoted by the epoch ID (eideid) of its last update. This eideid is utilized during the global commit phase (§3.4) to detect stale reads (i.e., cases where the local snapshot was superseded by concurrent remote updates).
Following execution, the transaction’s input, read-set, and write-set are stored in a $TxnRecord$ and appended to the current batch. Upon reaching capacity or after a timeout, the batch is asynchronously broadcast to all replicas ($BroadcastBatchAsync()$). Simultaneously, these batches are persisted into a linear, per-replica local log, ensuring availability for the subsequent commit protocol.
@@
Figure 3 illustrates the organization of locally executed transactions within the log of batches, where arrows denote data dependencies. In this example, $7$ transactions are distributed across two batches. Transactions $tx1$, $tx2$, $tx4$, and $tx7$ are independent; they operate on disjoint keys and read directly from the consistent snapshot. In contrast, $tx3$, $tx5$, and $tx6$ are dependent transactions, reading values written by uncommitted local operations from the temporary state. Specifically, $tx3$ depends on $tx1$ and $tx2$ (reading data item $A$ and writing to $B$), which implies an execution order where $tx3$ succeeds its predecessors. Similarly, $tx5$ depends on earlier writes. Transaction $tx6$, reading both $A$ and $B$, depends on both $tx5$ and $tx3$. The blue circles encircle these dependency groups, which we designate as Transaction Chains (TxnC). These chains serve as the atomic units for conflict detection during the global validation phase, as discussed in detail in Section 3.5.
@@ 3.3. Asynchronous Replication
11:    for $i\leftarrow PoA\_heads[self.rid]+1$ to $ack.bid$ do $\triangleright$ If a batch and its predecessors are confirmed available, broadcast the PoA until the latest available one.
@@
Figure 4. Replica A’s view of the logs: Brown means committed, blue means available and has PoA, green means broadcast but without PoA, and white means received but not available. The red dotted line is the coordinator’s consistent cut.
Upon receiving a batch from a peer, a replica appends the entry to the corresponding source replica’s log, marks it as broadcast, and sends an acknowledgment ($ReceivedBatch()$). The source replica aggregates these acknowledgments in $ReceivedAcknowledgement()$. After securing a quorum of $f+1$ acknowledgments, the batch transitions to the available state (Line 10). At this point, the source replica broadcasts a Proof of Availability (PoA), but only when all preceding batches in the sequence have already reached the available state and issued their own PoAs. This constraint ensures that availability is certified only for contiguous log prefixes. The PoA serves as a guarantee that the batch is persisted on at least one correct replica, satisfying the $f+1$ availability requirement.
The PoA generation process (Line 11) iteratively scans the log from the current PoA_headPoA\_head (the index of the last batch with a transmitted PoA) up to the newly available batch. Again, to ensure log contiguity, a PoA is broadcast for a batch only if all its predecessors are also marked available and the PoA_headPoA\_head is advanced. Upon receiving a PoA ($ReceivedPoA()$), peer replicas update the corresponding batch’s state to PoA_sentPoA\_sent. In addition, if a replica receives a PoA for a batch that is absent from its local log, it broadcasts a fetch request to retrieve the missing batch.
@@ 3.4. Global Epoch Commit
2: function StartConsensus $\triangleright$ At the coordinator
@@
4:    ProposeValue(PoA_headsPoA\_heads)
@@
7:   for all ridrid do
8:    for $i\leftarrow last\_committed[rid]$ to cut_bids[rid]cut\_bids[rid] do
@@
Periodically (every epoch), a designated coordinator replica (typically the Raft leader) initiates a consensus round triggered by a local timer ($StartConsensus()$). The coordinator proposes a ”consistent cut” of the global log, consisting of the vector of batch indices currently marked as PoA_sent in its local view (PoA_headsPoA\_heads). Standard consensus protocols, such as Raft (diegoRaft2014), are used for this proposal. Unlike systems that depend on synchronized physical clocks or centralized sequencers for ordering (thomsonCalvinFastDistributed2012; fanOceanVistaGossipbased2019; luEpochbasedCommitReplication2021; zhougeogauss2023), Minerva reaches agreement directly on the set of available batches. This approach establishes a consistent global checkpoint without assuming tight clock synchronization or imposing hard real-time constraints.
@@ 3.5. Validation and Deterministic Re-execution
Algorithm 4 Validation and Commit
1: var current_eid : int $\triangleright$ Index of the current epoch
@@
After collecting the batches_to_commitbatches\_to\_commit, replicas independently complete the validation and re-execution phase without further coordination.
We define two types of invalid transactions. The first type is stale transactions: those that read a data version from a consistent snapshot that has been superseded by a committed update prior to the current epoch. We use the epoch ID (eideid) as the write version to detect stale reads (Line 8 in Algorithm 5). In addition, we enforce cascading invalidation: any transaction that depends on a previously invalidated transaction, even across epochs, is recursively marked as stale. The second type is conflicting transactions, which include read-write or write-write conflicts in the current epoch.
@@
Following conflict resolution, the write-sets of all valid transactions are committed to durable storage. A key aspect of our algorithm is to deterministically re-execute the conflicting transactions instead of aborting them. To ensure serializability, stale and conflicting transactions are first arranged into a total order determined by their tidtid and the predefined priority of their source replica. This queue is then executed using a deterministic locking manager, derived from Calvin (thomsonCalvinFastDistributed2012), to enforce serializable execution according to the sorted order while maintaining concurrency. Deterministic locking requires prior knowledge of the read and write sets, which are available in our design because they are tracked during the initial OCC phase. Once the results of the re-executed transactions are persisted, the system responds to clients with the results.
@@ 3.6. Detecting Conflicting Transactions
Given a batch of transactions, we must efficiently identify write-write and read-write conflicts among them to ensure serializability (bernsteinconcurrency1981; ceriOCC1982). The conflict detection process must be deterministic, ensuring that every replica independently reaches the exact same commit decision; optimal, minimizing the number of re-executions; and computationally efficient, ensuring that the validation logic does not become a bottleneck in the commit critical path.
@@
25:   var $valid\_set\leftarrow$ SolveMWIS(CGCG)
26:   var $conflicting\_set\leftarrow CG.vertices\setminus valid\_set$
27:   return $(conflicting\_set\cup stale\_txns,\ valid\_set)$
Figure 5. Conflict graph of 6 transaction chains, arrows mean read-write or write-write conflicts, and the red box shows the conflicting transaction chains (TxnCTxnC) that are invalidated and to be re-executed.
Algorithm 5 shows the approach. We first construct the conflict graph ($CreateConflictGraph()$) through an iterative scan of all write-sets, linking transaction chains that have read-write or write-write contention on shared data items. Subsequently, the resulting graph is processed by the MWIS solver to identify the optimal subset of non-conflicting transactions. The chains that are excluded from the independent set are designated as conflicting. In cases where multiple optimal solutions exist, the solver may employ a deterministic tie-breaking rule, such as prioritizing the transaction chain with the lowest starting tidtid.
An example is shown in Figure 5, where $txnc2$ represents a heavy chain containing 5 transactions, while all other chains ($txnc1$, txnc3txnc3txnc6txnc6) each contain a single transaction. $txnc1$, $txnc2$, and $txnc3$ are in conflict because they may all have accessed the same data item. However, the MWIS solver should remove both $txnc1$ and $txnc3$ because $txnc2$ has the highest weight. For $txnc4$, $txnc5$, and $txnc6$, a possible scenario is that there are two data items, $a$ and $b$: $txnc4$ writes to $a$, $txnc5$ writes to both $a$ and $b$, and $txnc6$ reads from $b$. In this case, removing $txnc5$ is optimal.
Another issue to consider is cross-epoch and cross-batch dependent transactions. For cross-epoch dependencies, if a transaction depends on other transactions but they are not committed in the same epoch, then the transaction in the second batch must be derived from the previous consistent snapshot, so it is treated the same as a stale transaction. For dependencies across batches but still within the same epoch, the transactions are placed in the same global batch, and the links between the transaction chains across batches can be used to merge them into a single chain when constructing the conflict graph.
@@ 3.7. Correctness
We show that Minerva guarantees One-Copy Serializability (1SR) (bernsteinconcurrency1981) under crash failures. That is, all replicas observe the same serializable execution order of transactions. A transaction is defined as a series of read, write, or delete operations, where each operation accesses only a single data item. We do not assume any range queries or transactions with unbounded read/write sets, which can be supported with additional mechanisms such as index or partition versioning, or client-side locks (levandoskiMultiVersionRangeConcurrency2015).
@@
The coordinator only proposes the consistent cut with indices of batches that have received a PoA. Furthermore, a replica only broadcasts a PoA for a batch at index $i$ if all predecessor batches (indices <i<i) are already marked as available and have sent their PoAs (Line 11 of Algorithm 2). Therefore, by induction, if the coordinator observes a PoA for index $j$ (the cut index), then it is guaranteed that PoAs (and thus availability) exist for all indices $0,1,\dots,j-1$.
The consensus protocol (e.g., Raft) guarantees that the consistent cut indices are agreed upon by all correct replicas in the same order. Since any batches with indices less than the cut index are guaranteed to be available by Lemma 3.1, all these batches are guaranteed to be available. Because the batches are uniquely identified by monotonically increasing indices, their order will be identical across all correct replicas. ∎
@@
At each epoch, the same set of transactions is ordered to be executed in a serializable manner among all correct replicas.
@@
By Lemma 3.2, all correct replicas observe the same set of batches to commit, and thus the same set of all transactions $S_{all}$. The $CreateConflictGraph()$ function in Algorithm 5 is deterministic, as it detects all read-write and write-write conflicts among transactions (chains). The MWIS solver is also deterministic and returns the same optimal set of transactions $S_{opt}$ without read-write or write-write conflicts at each replica. Thus, there is no dependency among these transactions, and they can be considered serializable (thomasianDistributedOptimisticConcurrency1998).
The remaining transactions $S_{re}$ are re-executed using a deterministic locking mechanism. By the definition of deterministic locking (thomsonCalvinFastDistributed2012), $S_{re}$ is also deterministically serialized among all replicas.
Since $S_{re}$ is always executed after $S_{opt}$, and $S_{re}\cup S_{opt}=S_{all}$, the execution order of $S_{all}$ is serializable. ∎
@@
Base Case: Let $E_{0}$ be the initial epoch; it has no transactions and is trivially serializable.
@@
Induction Step: By the induction hypothesis, at the beginning of $E_{k}$, all replicas must start in the same state (the consistent snapshot). Thus, all transactions in epoch $E_{k}$ are validated and committed against the same consistent snapshot.
By Lemma 3.3, the commit procedure produces an execution order $S_{k}$ of all transactions in $E_{k}$ equivalent to a serial execution at all replicas, such that the state at the end of the epoch is $E_{k}=E_{k-1}+S_{k}$.
Since the state at the end of epoch $E_{k-1}$ is consistent among all replicas, and $S_{re}$ is serializable, the state at the end of epoch $E_{k}$ is also consistent among all replicas. Thus, $\sum_{i=1}^{k}S_{i}$ is $S_{ser}$. ∎
@@ 4. Implementation
We implement Minerva (minerva) as a distributed transactional key-value store. The system is developed in C# using .NET 9.0. For the network layer, we employ raw TCP sockets to handle all inter-replica and client interactions. We utilize MemoryPack (memorypack), a high-performance binary serializer for the communication protocol.
@@
Our deterministic re-execution engine implements a deterministic locking scheduler derived from Calvin (thomsonCalvinFastDistributed2012). Transactions are first assigned a global order based on a tuple of the source replica’s static priority and the transaction’s original ID (tidtid). A centralized lock manager enforces this schedule using per-key wait queues. Since the read and write sets are already known from the initial OCC attempt, lock acquisition is deterministic and deadlock-free. The lock manager maintains a dependency counter for each pending transaction; as active transactions complete and release their locks, the manager processes the relevant queues. When a transaction’s dependency counter reaches zero, it is scheduled for execution on a dedicated worker thread.
@@
As established in §3.5, we first need to identify all transaction chains within batches from the same replica. Using the local dependencies that are tracked during the OCC phase, the task is to find all connected components within the local dependency graph (distinct from the global conflict graph). This graph treats each transaction as a vertex, with edges linking transactions that share a local dependency. Then, a depth-first traversal partitions the graph into disjoint sets, producing the transaction chains.
Next, we use an open-source C++ graph library (Solver) that provides both exact and approximate solvers for the MWIS problem. For the exact solution, an integer linear programming (ILP) optimization method (wolseyIntegerProgramming2020) is used. Although this approach incurs higher computational latency, it guarantees the identification of the optimal independent set and minimizes the volume of transactions to re-execute. The approximate algorithm is based on a popular greedy heuristic (sakaigreedymwis2003), which prioritizes decision speed over optimality. The performance trade-offs between these two approaches are analyzed in Section 5.5.
@@
Minerva eliminates the need for the blocking 2PC process to ensure atomicity within the replication cluster because of the replication protocol and deterministic validation. Replicas can thus independently converge on the same commit or abort decisions for every transaction without additional coordination.
The Raft cluster is also used as the configuration manager. The Raft leader assumes the role of the coordinator replica, which is responsible for initiating the periodic global commit process. This also effectively handles coordinator failures, as Raft guarantees that a new leader will be elected to resume the commit process. We use an open-source implementation from the .Next library (DotNext).
Finally, we propose the following potential recovery mechanisms to handle replica failures. While failed replicas can recover by fetching the missing batches from other replicas based on the global log, the global log should be persisted on the disk with a write-ahead log, allowing the recovered replica to only gather the missing batches without the need to re-execute all transactions within the epoch. Furthermore, storage nodes can employ a localized primary-backup mechanism with asynchronous backup. Since transactions are always written to the database snapshot first, we do not need to consider the consistency problem with the backups by using epoch as versioning.
@@
We implement an adaptive high-contention mode to reduce performance impact under heavily contended workloads. This mode is triggered when the percentage of re-executed transactions exceeds a configurable threshold (default $50\%$) for a sustained duration (default $200$ ms). In this mode, replicas suspend the optimistic execution phase entirely. Instead, incoming transactions are statically probed to extract their read and write sets and are immediately forwarded to the current batch for replication. During the global commit phase, these transactions bypass the conflict detection phase and are automatically routed to the deterministic execution queue.
@@
To prevent unbounded storage growth, obsolete log entries are garbage collected when they are no longer needed, that is after they have been fully processed by the entire cluster. To track this, replicas periodically broadcast their progress (their latest committed epoch). Each node maintains a local view of the cluster’s ”low watermark,” which is the highest epoch index that has been universally committed. Once this watermark advances beyond epoch $i$, replicas safely truncate the log up to that point. In addition, the local temp_statetemp\_state is cleared at each epoch boundary, effectively garbage collecting the invalid transactions.